Marian Plunket owns her own business and is considering an investment. If she undertakes the​ investment, it will pay $ 5 comma 880 at the end of each of the next 3 years. The opportunity requires an initial investment of $ 1 comma 470 plus an additional investment at the end of the second year of $ 7 comma 350. What is the NPV of this opportunity if the interest rate is 1.6 % per​ year? Should Marian take​ it?
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=5880/1.016+5880/1.016^2+5880/1.016^3

=17090.22

Present value of outflows=1470+7350/1.016^2

=8590.33

NPV=Present value of inflows-Present value of outflows

=17090.22-8590.33

=$8499.89(Approx).

Hence since NPV is positive;investment must be taken.